For the function f (x)= 1-(1)/(x+2)^(2),find the following:

(i)f(-x) (ii)f(x +k)


1
Expert's answer
2022-03-15T05:18:36-0400

The function has the following form: "f(x)=1-\frac{1}{(x+2)^2}".

(i) Substitute "-x" into the expression for function "f(x):" "f(-x)=1-\frac{1}{(2-x)^2}";

(ii) Substitute "x+k" into the expression for the function "f(x):" "f(x+k)=1-\frac{1}{(x+k+2)^2}"